If let z= x+ iy , |z|= 8 then value of |(conj(z)+48)/(3z + 4)|=.....
( 2 , 4 , 8 , 16 )

Found 2 solutions by Edwin McCravy, math_tutor2020:
Answer by Edwin McCravy(20056)

If let z= x+ iy , |z|= 8 then value of |(conj(z)+48)/(3z + 4)|=.....
( 2 , 4 , 8 , 16 )
This is a multiple-choice problem with 4 potential answers.

So, it must be true that the answer is the same no matter what x and y are,

as long as |z| = |x+iy| = 8

That's because your teacher couldn't have given you only 4 numerical choices if
there could be different values for different values of x and y.
 
So, instead of bothering to prove it for all values of x and y, just choose this
easy case:

x=8, y=0, then z = 8+0i = 8-0i = conj(8+0i) = 8, then

|(conj(z)+48)/(3z + 4)| = |8+48)/[(3)(8) + 4]| = |56/28| = |2| = 2.

So the answer is 2. 

Note: if your problem had been stated this way: 

if |z| = 8 then prove that the value of |(conj(z)+48)/(3z + 4)| = 2,

then there would be a lot more work.  So be on the lookout for multiple
choice questions like this with variables which have no variables in the
choices.  For if you run into such a problem, just substitute easy numbers
for the variables and whatever you get will bound to be the correct answer.
You'll save lots of time.

Edwin

Answer by math_tutor2020(3817)


z*conj(z) = |z|^2
z*conj(z) = 8^2
z*conj(z) = 64
conj(z) = 64/z

Apply substitution and simplify.
(conj(z) + 48)/(3z+4)
= (64/z + 48)/(3z+4)
= (64 + 48z)/(z(3z+4)) ... multiply top and bottom by z
= (16(3z+4))/(z(3z+4))
= 16/z
We have something much more simple to work with.

Which means,
| (conj(z) + 48)/(3z+4) |
= | 16/z |
= |16| / |z|
= 16/8
= 2

Therefore, | (conj(z) + 48)/(3z+4) | = 2 when |z| = 8